Menlo Company distributes a single product. The company’s sales and expenses for last month follow:

Total    Per Unit
  Sales $ 302,000 $ 20     
  Variable expenses 211,400 14     
  Contribution margin 90,600 $ 6     
  Fixed expenses 72,000
  Net operating income $   18,600

Required:

1. What is the monthly break-even point in unit sales and in dollar sales?

2. Without resorting to computations, what is the total contribution margin at the break-even point?

3-a. How many units would have to be sold each month to earn a target profit of $41,400? Use the formula method.

3-b. Verify your answer by preparing a contribution format income statement at the target sales level.

4. Refer to part 3 and now assume that the tax rate is 30%. How many units would need to be sold each month to an after-tax target profit of $41,400?  (Round the final answer to the nearest whole number.)

5. Refer to the original data. Compute the company's margin of safety in both dollar and percentage terms. (Round your percentage answer to 2 decimal places (i.e .1234 should be entered as 12.34).)

6. What is the company’s CM ratio? If monthly sales increase by $67,000 and there is no change in fixed expenses, by how much would you expect monthly net operating income to increase?
Answer #1

1. Breakeven point in unit sales = $ 72,000 / $ 6 = 12,000 units

Breakeven point in dollars = $ 72,000 / 0.30 = $ 240,000

2. $ 72,000.

3-a. Units that would have to be sold to earn a target profit of $ 41,400 = $ ( 72,000 + 41,400) / $ 6 = 18,900 units.

3-b.

Sales ( 18,900 x $ 20) $ 378,000
Variable Costs 264,600
Contribution Margin 113,400
Fixed Costs 72,000
Net Income $ 41,400

4. Before tax target profit = $ 41,400 / ( 1 - 0.30 ) = $ 59,143

Sales in units required to earn after tax profit of $ 41,400 = $ ( 72,000 + 59,143 ) / $ 6 = 21,857.17 units

5. Margin of Safety = $ 302,000 - $ 240,000 = $ 62,000

Margin of Safety % = $ ( 302,000 - 240,000 ) / $ 302,000 * 100 = 20.53 %

6. CM ratio = $ 6 / $ 20 * 100 = 30 %

If monthly sales increase by $ 67,000, monthly net operating income will increase by $ 67,000 x 30 % = $ 20,100.